A quantity of gas has a volume of 0.20 cubic meter and an absolute temperature of 333 degrees kelvin. When the temperature of the gas is raised to 533 degrees kelvin, what is the new volume of the gas? (Assume that there's no change in pressure.)
A. 0.0006 m3
B. 0.2146 m3
C. 0.2333 m3
D. 0.3198 m3


Answer by Alan3354(69443)

PV = nRT
If only V and T vary:
V = kT
0.2 = k*333
k = 0.2/333
V at 533 = k*533
V = 0.2*(533/333) =~ 0.3201 cubic meters